# A Fully Automated Pipeline for Conversational Discourse Annotation: Tree Scheme Generation and Labeling with Large Language Models

### v1

@@ 6 Open-Source Models for Annotation
Table 4 presents the results of using two open-source models, Mistral-7B-Instruct-v0.3888https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct, [^9] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct compared to the closed-source models GPT-3.5 and GPT-4o for the annotation steps with two approaches, frequency-guided and non-frequency-guided optimal split selection with backtracking, which were selected based on their strong performance with other models.
The results indicate that despite having fewer parameters than the closed-source models, both open-source models significantly outperform GPT-3.5. Specifically, when using the non-frequency-guided approach, both open-source models achieve performance close to GPT-4o while markedly surpassing GPT-3.5. However, with the frequency-guided approach, Llama continues to outperform GPT-3.5 significantly, but both models’ performance metrics fall short of GPT-4o. This trend underscores the substantial performance gains achieved by combining the frequency-guided approach with GPT-4o, as discussed in Section 5.4.2.
@@ 7 Evaluation on Switchboard Dialog Act Corpus
Table 6: Evaluation of dialog act annotations from the SWBD-DAMSL annotation scheme, generated by GPT-4o, on a randomly selected set of 260 utterances from the Switchboard Dialogue Act Corpus using non-binary open-ended questions tree, and a frequency-guided optimal split selection tree.
@@ 8 Conclusions
Potential directions for future research, motivated by the current limitations, include: (1) exploring larger open-source models for improved taxonomy generation and annotation; (2) conducting experiments on other domains, such as classroom discourse and task-oriented dialog systems; (3) incorporating human feedback to allow the model to self-correct and improve annotation accuracy; and (4) enabling self-refinement of the taxonomy by adapting to new, previously unseen dialog examples.
@@
Dialog data often contains personal or sensitive information, making it essential to anonymize and handle data securely when applying the proposed approach to individual datasets. This is crucial to safeguarding privacy rights. Beyond this consideration, we do not anticipate any significant risks associated with this work or the use of the proposed framework.
@@ Appendix I Prompt for Constructing a Yes/No Questions Tree
Figure 12: The prompt is designed to create a decision tree that enables splitting data into more than two groups and that specifically instructs the LLM to form one group that exclusively contains the single most frequent class

### v2

@@ 7 Open-Source Models for Annotation
Table 4 also presents the results of using two open-source models, Mistral-7B-Instruct-v0.3555https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct, [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct compared to the closed-source models GPT-3.5 and GPT-4o for the annotation steps with two approaches, frequency-guided and non-frequency-guided optimal split selection with backtracking, which were selected based on their strong performance with other models.
These results indicate that despite having fewer parameters than the closed-source models, both open-source models notably outperform GPT-3.5. Specifically, when using the non-frequency-guided approach, both open-source models achieve performance close to GPT-4o while markedly surpassing GPT-3.5. However, with the frequency-guided approach, Llama continues to outperform GPT-3.5 noticeably, but both models’ performance metrics fall short of GPT-4o. This trend underscores the substantial performance gains achieved by combining the frequency-guided approach with GPT-4o, as discussed in Section 6.2.
@@ 8 Evaluation on Switchboard Dialog Act Corpus
Table 8: Evaluation of dialog act annotations from the SWBD-DAMSL annotation scheme, generated by GPT-4o, on a randomly selected set of 260 utterances from the Switchboard Dialogue Act Corpus using non-binary open-ended questions tree and a frequency-guided optimal split selection tree.
@@ 9 Conclusions
Potential directions for future research, motivated by the current limitations, include: (1) exploring larger open-source models for improved taxonomy generation and annotation; (2) conducting experiments on other domains, such as classroom discourse and task-oriented dialog systems; (3) incorporating human feedback to allow the model to self-correct and improve annotation accuracy; (4) enabling self-refinement of the taxonomy by adapting to new, previously unseen dialog examples; and (5) allowing the annotation step to select multiple candidate branches/labels, followed by a final evaluation step that explicitly compares the utterance against the chosen candidates’ class definitions to determine the most suitable class.
@@
Dialog data often contains personal or sensitive information, making it essential to anonymize and handle data securely when applying the proposed approach to individual datasets. This is crucial for protecting privacy rights. Beyond this consideration, we do not anticipate any significant risks associated with this work or the use of the proposed framework.
@@ Appendix I Prompt for Constructing a Yes/No Questions Tree
Figure 12: The prompt is designed to create a decision tree that enables splitting data into more than two groups and that specifically instructs the LLM to form one group that exclusively contains the single most frequent class.